Exercise 2. If x > 0, show that there exists n ∈ N such that 1/n < x.


1
Expert's answer
2021-01-14T20:02:24-0500

If 1/n >= x > 0 for all n ∈ N, then for all n ∈ N 1/x > n.

Let's prove that such real number doesn't exist.

Let A be a set of real numbers "A = \{a\in \mathbb{R}: \exist n\in \mathbb{N}\,\, a<n\}" ,

"B=\mathbb{R}\setminus A = \{b\in \mathbb{R}: \forall n\in \mathbb{N}\,\, b\geq n\}" .

"\mathbb{N}\subset A" implies "A\ne \empty". If, in addition, "B\ne \empty", then "\forall a\in A \,\forall b\in B\, a< b" .

Indeed, there exists "n\in N", such that a<n and, by the definition, "n\leq b" , therefore a<b.

So we have Dedekind's section of the set of real numbers.

By the Dedekind's axiom of completeness there exists "c\in \mathbb{R}" , such that "\forall a\in A \,\forall b\in B\, a\leq c\leq b" .

We have "c-1<c", therefore, "c-1\in A".

We have "c+1>c", therefore, "c+1\in B".

By the definition of the set A we have "\exist n\in \mathbb{N} c-1<n" and, hence, c+1 < n+2. The last inequality means that c+1 must belong to A. This contradiction proves that "B=\empty" .